OpenStudy (anonymous):

4.Write the equation of the line which passes through (2, –3) and is perpendicular to y = 4x + 7 in standard form. (2 point) plz helpp

OpenStudy (anonymous):

What is the slope of the given line y = 4x + 7?

OpenStudy (anonymous):

? * ?

hero (hero):

y - y1 = m(x-x1) y + 3 =-1/4(x - 2) y + 3 = -x/4 + 1/2 y = -x/4 + 1/2 - 3 y = -x/4 + 1/2 - 6/2 y =-x/4 -5/2

OpenStudy (anonymous):

thanks

hero (hero):

yw :D